Circular No. 7295

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1999 U3

R. Huber reports the discovery by LINEAR of a comet. Additional observations have been reported following posting on the NEO Confirmation Page:

1999 UT α_{2000} δ_{2000} m_{1} Observer
Oct. 30.32373 3^{h}27^{m}14.^{s}49 +58^{o}17′44′′.3 18.5 LINEAR
30.34009 3 27 14.31 +58 17 50.7 19.1  _{”}
30.37274 3 27 14.12 +58 18 01.6 19.6  _{”}
30.38902 3 27 13.82 +58 18 06.5 18.6  _{”}
31.33285 3 27 07.79 +58 22 31.9 17.7 Garradd
31.33578 3 27 07.79 +58 22 32.6 17.4  _{”}
31.33840 3 27 07.73 +58 22 33.3  _{”}
31.34194 3 27 07.71 +58 22 34.2 17.8  _{”}
31.71834 3 27 05.36 +58 24 04.7  Šarounová
31.72825 3 27 05.26 +58 24 07.2 17.0  _{”}
31.73023 3 27 05.25 +58 24 07.7 16.2  _{”}
31.74209 3 27 05.19 +58 24 11.5 17.2 Tichá
31.74374 3 27 05.20 +58 24 11.8  _{”}
31.74492 3 27 05.16 +58 24 11.9  _{”}
31.74605 3 27 05.18 +58 24 12.5  _{”}
31.74838 3 27 05.15 +58 24 13.1  _{”}
31.75147 3 27 05.11 +58 24 14.1  _{”}
31.82326 3 27 04.36 +58 24 33.4  Busch
31.83704 3 27 04.29 +58 24 37.7  _{”}
31.88041 3 27 03.74 +58 24 49.7  Šarounová

LINEAR (Lincoln Laboratory ETS, Socorro). Observers M. Bezpalko, M. Blythe, F. Shelly, M. Elowitz, and R. Huber. Measurers J. Stuart, H. Viggh, and R. Sayer. 1.0-m f/2.15 reflector + CCD. Short tail.

G. Garradd (Oracle, AZ). Biosphere2 Observatory 0.5-m f/8.1 reflector + CCD. Diffuse coma of diameter ∼ 10′′ and faint tail extending ∼ 15′′ to the southwest.

L. Šarounová (Ondřejov). 0.65-m f/3.6 reflector + CCD. Condensed nu-

cleus; faint tail 25′′ long in p.a. 215^{o}. J. Tichá and M. Tichý (Kleť). 0.57-m f/5.2 reflector + CCD. Slightly diffuse

with 14′′ tail in p.a. ∼ 215^{o}. M. Busch, R. Kresken, and R. Stoss (Heppenheim, Germany). 0.45-m f/4.4

reflector + CCD. Tail > 18′′ long in p.a. 225^{o}.

1999 October 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green